OpenStudy (anonymous):

how do i label a unit circle that is divided into 16 slices

OpenStudy (anonymous):

start from the x-line

OpenStudy (anonymous):

where theta = 0

OpenStudy (anonymous):

now uve got to make 16 slices whts the total angle of the circle?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

see the total angle of a circle is 360 degrees now uve got to divide into 16 parts 360/16 = 45/2 = 22.5 degrees

OpenStudy (anonymous):

so now u have to label at intervals of 22.5 degrees..so now mark the multiples of 22.5

OpenStudy (anonymous):

u understand?

OpenStudy (anonymous):

yes I understand that part. Now, how do I find the roots of x^16-1.

OpenStudy (anonymous):

is the unit circle to be divided into 16 evenly? if not you could just have one chunk as 346 degrees then the last 15 as the other fifteen divisions

OpenStudy (anonymous):

ok so now uve got to mark the roots of x^16 = 1

OpenStudy (anonymous):

weve divided the circle into 16 parts...so now we just mark all the multiples of 22.5 those are the roots of x^16 = 1

OpenStudy (anonymous):

The question is basically ask to find roots of x^16-1 then it ask what information does it contain about the trigonometric functions. So I am not sure if it would be even parts or not. I figured dividing the circle would be my easiest approach but doesn't look so easy

OpenStudy (anonymous):

okay, give me one sec I am making the circle

OpenStudy (anonymous):

see for the trig part the roots turn up as coordinate points, the values of which are the values of different trig ratios

OpenStudy (anonymous):

joemath please dont strt using the euler form

OpenStudy (anonymous):

>.> fine haha

OpenStudy (anonymous):

cz she has no idea right now

OpenStudy (anonymous):

Okay once i have the circle what do I do

OpenStudy (anonymous):

Geeze Him, way to make me feel better about myself. lol He is right I am clueless

OpenStudy (anonymous):

no i didnt mean it that way...its just that u havent studied it maybe...thats y..no offence..if u want i cn teach u

OpenStudy (anonymous):

him's judgement is correct, i would have complicated things a lot more by putting my solution up

OpenStudy (anonymous):

No I am not offended. Its new to me. Yes its hard and confusing but i enjoy it and want to learn it. My teachers so far have not been so good teaching it

OpenStudy (anonymous):

nvm..we can be ur teachers..u chose to ignore me last night...bt i still wld like to teach u

OpenStudy (anonymous):

so u have the unit circle..and now we can start marking the points

OpenStudy (anonymous):

yes I have the circle and points

OpenStudy (anonymous):

make a mark on the circle where theta=0

OpenStudy (anonymous):

have u marked all the points?

OpenStudy (anonymous):

yes so at 0 you want me to put theta

OpenStudy (anonymous):

no no...if uve marked all the points...u can see these points represent the roots of ur eqn

OpenStudy (anonymous):

come to chat..its much faster

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

group chat on the bottom right is flashing orange

OpenStudy (anonymous):

u understand now? look at the figure...(a,b) is ur complex number, now tell me whts cos theta?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Hey liz m sorry m not going to be there this evening....please manage on ur own tonyt..im sure loads f ppl will be there fr u...dont worry....ill be there at night..sorry